Word Problem: Can someone please tell me if I have the right answer?
During rush hour, Fernando can drive 20 miles using the side roads in the same time that it takes to travel 15 miles on the freeway. If Fernando's rate on the side roads is 9 mi/h faster than his rate on the freeway, find his rate on the side roads.
Choices are:
A) 38 B) 36 C) 27 D) 29
The answer I get is C....

Answer by checkley71(8403)

D=R*T DISTANCE = RATE * TIME OR T=D/R & SEEING AS THE TIMES ARE THE SAME WE SET EACH TRIP EQUAL THUS 20/(x+9)=15/x or 20x=15x+135 or 5x=135 or x=135/5 or x=27 or x+9=36 for the speed on the side roads.
ANS. IS B)36 MPH